OpenStudy (anonymous):

cos(5pie/3)= A. root 3/2 B. root 2/2 C. -root 2/2 D. 1/2

OpenStudy (solomonzelman):

cos(5π/3) = cos(6π/3 - π/3) = cos(6π - π/3) = cos(2π)cos(π/3) + sin(2π)sin(π/3) = 1 • cos(π/3) + 0 • sin(2π/3) = cos(π/3)

OpenStudy (anonymous):

my cos says root 1/2

OpenStudy (solomonzelman):

cos(5π/3) = cos(6π/3 - π/3) = cos(6π - π/3) = cos(2π)cos(π/3) + sin(2π)sin(π/3) = 1 • cos(π/3) + 0 • sin(π/3) = cos(π/3)

OpenStudy (anonymous):

so D?

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} \cos \left( {\frac{{5\pi }}{3}} \right) = \cos \left( {\pi + \frac{{2\pi }}{3}} \right) = \hfill \\ \hfill \\ = \cos \pi \cos \left( {\frac{{2\pi }}{3}} \right) - \sin \pi \sin \left( {\frac{{2\pi }}{3}} \right) = ... \hfill \\ \end{gathered} \]

OpenStudy (solomonzelman):



OpenStudy (anonymous):

is it D?

OpenStudy (solomonzelman):

yes, D is right

OpenStudy (michele_laino):

hint: \[\begin{gathered} \cos \left( {\frac{{5\pi }}{3}} \right) = \cos \left( {\pi + \frac{{2\pi }}{3}} \right) = \hfill \\ \hfill \\ = \cos \pi \cos \left( {\frac{{2\pi }}{3}} \right) - \sin \pi \sin \left( {\frac{{2\pi }}{3}} \right) = ... \hfill \\ \end{gathered} \]

OpenStudy (solomonzelman):

you can do it as (2π-π/3), it seems easier this way a bit.... (to me at least)

OpenStudy (anonymous):

Thank you

OpenStudy (solomonzelman):

yw

OpenStudy (solomonzelman):

and I still had a typo, once agian

OpenStudy (solomonzelman):

cos(5π/3) = cos(6π/3 - π/3) = cos(2π - π/3) ^ I said that was 6π (when it is 2π) = cos(2π)cos(π/3) + sin(2π)sin(π/3) = 1 • cos(π/3) + 0 • sin(2π/3) = cos(π/3)

OpenStudy (solomonzelman):

in any case, if you have questions ask, if not then good luck